# Almost commuting self-adjoint operators and measurements

### v2

@@ 1 Introduction
In the First International Congress of Basic Science held in 2023, David Mumford gave the opening plenary lecture on Consciousness, robots and DNA ([32]). At the very beginning, he cautioned the audience that there would be no mathematics in his talk. Nevertheless, by the very end of this magnificent lecture, when he ended his talk on DNA being a measuring instrument opening a Pandoraβs box, he proposed a problem in -algebras about non-commuting self-adjoint operators to be approximated by commuting ones.
@@
| (eβ1.2) |
In the case that and is any finite dimensional Hilbert space (no bound on the dimension) the exactly the same problem is known as von-Neumann-Kadison-Halmos problem for almost commuting self-adjoint matrices (see [11] and [12]). In that case, the problem has an affirmative solution [20] (see also [8] and [13], and, for a fixed dimension, see [31]). In infinite dimensional Hilbert space, as observed by Mumford in his new book (p.182), the answer in general is negative (see also Example 4.6 of [8]). The affirmative solution for almost commuting self-adjoint matrices led the further study in the subject of weakly semi-projectivity in -algebras. It also has a deep impact on the Elliott program of classification of amenable -algebras (see [7], [24], [26] and [28]). The Mumfordβs problem opens a new door to the future applications of this part of -algebraβ theory in previous unexpected areas such as DNA and artificial intelligence ([33]).
@@
Let and There exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and are bounded self-adjoint linear operators on with () such that
| (eβ1.3) |
@@
Here and are the sets of -synthetic-spectrum, and essential -synthetic-spectrum, respectively (see Definition 2.7). In (eβ1.3), is the Hausdorff distance between the two compact subsets of the -dimensional unit cube. In the special case that is full, i.e., it contains the whole then the second part of (eβ1.3) is automatically satisfied.
@@
More generally, let be a Hilbert (right) module over a -algebraβ be the -algebraβ of all bounded module maps on with adjoints and let be the linear span of rank one module maps of the form (for all ), where Denote by the closure of which is a -algebra. We also have the following:
@@
Let and There exists satisfying the following: Suppose that is a countably generated Hilbert-module over a -unital purely infinite simple -algebraβ and are self-adjoint bounded module maps with () such that
| (eβ1.5) |
@@
(ii) is a countably generated Hilbert module over a -unital purely infinite simple -algebraβ
@@
-algebraβ of stable rank one such that has real rank zero and is simple.
@@
The second condition in (eβ1.7) means that a Fredholm index vanishes (see Definition 2.4). Note that both case (ii) and case (iii) are generalization of case (i). In case (ii), is stable (whenever is not a finitely generated projective Hilbert module), and is simple (like the Calkin algebra). In case (ii), we keep the feature that is simple and has real rank zero.
Case (i) in Theorem 1.3 is a more definite result than that of Theorem 1.1 in the case We actually prove a slightly more general result (see Theorem 5.3 below). One should also compare (i) in Theorem 1.3 to Theorem 1.1 of [17] (see Remark 5.6). In the case that however, the normal elements may not be approximated by normal elements in with finite spectrum. We also show that condition of the vanishing index associated with the pair of operators is necessary. Since it is well-known that spectrum (and the essential spectrum) of a bounded operator is notoriously unstable regarding perturbation, it is necessary to consider some version of -spectrum (which is usually larger than the spectrum). For we introduce the notion of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebraβ (see Definition 2.7). We will also show that, unfortunately, the version of Mumfordβs problem formulated above has a negative solution in general even the associated Fredholm index vanishes.
In quantum mechanics, one often heard βwhen the commutators go to zero, we recover the classical systemβ. If βrecoverβ means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [33] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all show that, mathematically, there are serious (topological) obstacles prevent this type of statements from being true. Nevertheless Theorem 1.1 with some interpretation in Remark 6.6 may provide an alternative statement.
The paper is organized as follows. Section 2 contains some notations used in the paper and gives the definition of -synthetic-spectrum and -near-spectrum for -tuples of self-adjoint operators, and discusses the relationship between them (see also Proposition 2.16). Section 3 is a study of -tuples of almost commuting self-adjoint elements in a unital purely infinite simple -algebra. Section 4 presents the proof of Theorem 1.1 and Theorem 1.2. In Section 5, we will discuss the case of a pair of almost commuting self-adjoint bounded module maps and provide a proof of Theorem 1.3. In section 6, the last section, we show why the Fredholm index does not appear in the statement of Theorem 1.1. We then present an example that an -tuple version of Theorem 1.3 does not hold in general. A final remark was added which contains the justification of the second condition in (eβ1.3) and its connection to the Mumford original question.
@@ 2 Preliminaries
Let be a -algebra. Denote by the set of all self-adjoint elements in Denote by the multiplier algebra (the idealizer of in ). Denote by and the group of invertible elements and the path connected component of containing the identity of respectively.
If and we write
| (eβ2.9) |
@@
Let be an infinite dimensional separable Hilbert space. Denote by the -algebraβ of all bounded linear operators on and by the -algebraβ of all compact operators on Denote by the quotient map.
@@
Let be a -algebraβ and be a Hilbert (right) module over (see [15]), or Hilbert -module. Denote by the -algebraβ of all bounded module maps on with adjoints. A rank one module map is a bounded module map of the form for all (and fixed ), where is the -value inner product on Denote by the linear span of rank one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
@@
Let be a unital -algebraβ and be an invertible element. Define Note that is a unitary. Define
| (eβ2.10) |
Then if and only if Let be such that is invertible in Then
| (eβ2.11) |
@@
Let In what follows, for each we choose a fixed integer such that Denote We write Then is -dense in Moreover, if
Let be a unital -algebraβ and be an -tuple of self-adjoint elements in with (). Fix Let be such that if if and is linear in and in Define, for
| (eβ2.12) | |||||
| (eβ2.13) |
Note that we do not assume that mutually commute and the product in (eβ2.13) has a fixed order.
@@
Let be a unital -algebraβ and for some Let us assume that Fix Suppose that is a compact subset and is a unital c.p.c. map such that
| (eβ2.19) | |||||
| (eβ2.20) | |||||
| (eβ2.21) |
@@
We often use the -near spectrum as it associated with a c.p.c. map, whenever the -tuple sufficiently almost commutative as shown in the next proposition (2.11). But it is not unique and depends on the choice of the c.p.c. map. On the other hand, the -synthetic-spectrum is always defined and in the computation world, it is computable. Their relationship is described in Proposition 2.15 below (see also Proposition 2.16). In practice, it is quite possible that it is more difficult to compute the spectrum of an operator where than to compute the norm of
@@
Fix For any there exists satisfying the following: Suppose that is a unital -algebraβ and with such that
| (eβ2.22) |
@@
Consider and quotient map Let Then Let be the -subalgebraβ of generated by Then for some non-empty compact subset of Let be the monomorphism given by with
@@
Then, for any and any finite subset there exists satisfying the following: Suppose that is a unital -algebraβ and are unital c.p.c. maps such that
| (eβ2.29) |
@@
| (eβ2.30) |
Moreover, if and then, for any and and finite subset we may also require that, for all
| (eβ2.31) |
@@
Put Let the quotient map. Then, since is a generating set of by (eβ2.33), is a unital homomorphism, and by (eβ2.32), Hence
| (eβ2.36) |
@@
| (eβ2.37) |
This leads a contradiction to (eβ2.34). The first part of the lemma then follows.
@@
| (eβ2.39) |
Another contradiction. So the lemma follows. β
@@
Let be as defined in 2.7, where (). There is satisfying the following: for any unital -algebraβ and any pair of self-adjoint elements with if then
| (eβ2.42) |
@@
Let and Choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebraβ ), that
| (eβ2.43) |
@@
| (eβ2.45) |
for any -tuple of self-adjoint elements in with (), by the choice of (and ) and by the second part of (eβ2.40) and (eβ2.42), we have, for each
| (eβ2.46) |
@@
| (eβ2.48) |
It follows that (keeping in mind of (eβ2.45))
| (eβ2.49) |
Combining (eβ2.49) with (eβ2.47), we obtain
| (eβ2.50) |
We conclude that, by (eβ2.50) and (eβ2.46),
| (eβ2.51) | |||
| (eβ2.52) | |||
| (eβ2.53) |
β
@@
(1) Suppose that is a unital -algebraβ and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of and then
| (eβ2.55) |
@@
| (eβ2.57) | |||
| (eβ2.58) | |||
| (eβ2.59) |
for any with for all whenever (if ) or (if ). Let We may assume that Then By (eβ2.59),
| (eβ2.60) |
By (eβ2.58) and the choice of applying Lemma 2.12,
| (eβ2.61) |
@@
| (eβ2.67) |
Next, let By (eβ2.64),
| (eβ2.68) |
Hence Thus
| (eβ2.69) |
It follows that
| (eβ2.70) |
Hence
| (eβ2.71) |
@@
| (eβ2.72) | |||
| (eβ2.73) | |||
| (eβ2.74) |
for any with if for some Choose We may assume that Hence It follows from (eβ2.74) that
| (eβ2.75) |
On the other hand, by the choice of (eβ2.58), (eβ2.73), and applying Lemma 2.12, we have
| (eβ2.76) |
@@
Let There exists satisfying the following: Suppose that is a unital -algebraβ and such that (), and
| (eβ2.79) |
@@
Let and be the quotient map. Put Then, by (eβ2.80), is normal. Let Then defined by for is a unital monomorphism. By the Choi-Effros Lifting Theorem ([4]), there is a sequence of unital c.p.c. maps such that Let be as in Proposition 2.15. Then, since is a homomorphism, for all large is a -near-spectrum for Applying Proposition 2.15, we may assume that, for all
| (eβ2.83) |
Note that, by (eβ2.82) and (eβ2.83), Let Note that Hence there are such that ()
| (eβ2.84) |
@@
| (eβ2.85) |
This is a contradiction. So the lemma follows. β
@@ 3 Purely infinite simple -algebras
Let be a compact subset of Define to be the union of all even -bricks (in with corners at ) which has a non-empty intersection with
@@
Let be a -algebraβ and a homomorphism. Denote by the restriction.
@@
Let be a compact metric space with property (F). Suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphismβ such that
| (eβ3.93) |
Then, for any and any finite subset there exists a set of finitely many mutually orthogonal projections with and such that
| (eβ3.94) |
@@
Recall that is finitely generated (). It follows that, for any unital -algebraβ
| (eβ3.95) |
Define by for all Then one computes that, by (eβ3.93),
| (eβ3.96) |
@@
Recall that, by S. Zhangβs theorem ([38]), has real rank zero. Since is injective, by applying Lemma 4.1 of [7], we obtain a set of mutually orthogonal non-zero projections with such that
| (eβ3.99) |
@@
| (eβ3.103) |
Define for all Then is a unital homomorphismβ from to with finite dimensional range and, by (eβ3.103), (eβ3.100) and (eβ3.98), for all
Finally, since in as we arranged earlier, by the definition of and the βMoreoverβ part of the statement also holds. β
@@
Let be a compact subset of Suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphism. If is any compact subset with property (F) such that where then there is a a unital injective homomorphismβ such that
| (eβ3.105) |
@@
| (eβ3.106) |
Fix any Suppose that is a compact subset of such that Since is compact, there is with such that
| (eβ3.107) |
@@
| (eβ3.109) |
Applying Theorem 3.5, there is a unital homomorphismβ with finite dimensional range such that
| (eβ3.110) |
@@
| (eβ3.115) |
Then, combining with (eβ3.110) and (eβ3.112), for all
| (eβ3.116) |
@@
| (eβ3.117) |
Then is a unital injective homomorphism. Moreover, for all by (eβ3.109), the definition of above and by (eβ3.116),
| (eβ3.118) | |||||
To see the second formula in (eβ3.105), we may write, from (eβ3.117), since factors through
| (eβ3.119) |
β
@@
Suppose that is a compact subset of and suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphism. If is any compact subset with property (F) such that where then there is a unital injective homomorphismβ such that
| (eβ3.120) |
@@
Let be a compact metric space which is a a compact subset of (for integer ) and be a unital purely infinite simple -algebra. Suppose that is a unital injective homomorphism. Then, for any and any finite subset there exists such that, for any integer with there exists a unital injective homomorphismβ such that
| (eβ3.122) |
@@
Let and There is satisfying the following: If is a unital purely infinite simple -algebraβ and with () such that
| (eβ3.123) |
then there is integer with and and a -brick combination and a unital injective homomorphismβ such that
| (eβ3.124) |
@@
| (eβ3.125) |
Let be the commutative -subalgebraβ generated by Then there exists a unital homomorphismβ such that Hence for some compact subset of Denote the isomorphism give above. We may assume that
Put Choose any with and By Theorem 3.6, we obtain a -brick combination and a unital injective homomorphismβ such that
| (eβ3.126) |

### v3

@@ 1 Introduction
In the First International Congress of Basic Science held in 2023, David Mumford gave the opening plenary lecture on Consciousness, robots and DNA ([33]). At the very beginning, he cautioned the audience that there would be no mathematics in his talk. Nevertheless, by the very end of this magnificent lecture, when he ended his talk on DNA being a measuring instrument opening a Pandora’s box, he proposed a problem in -algebras about non-commuting self-adjoint operators to be approximated by commuting ones.
@@
| (e 1.2) |
In the case that and is any finite dimensional Hilbert space (no bound on the dimension) the exactly the same problem is known as von-Neumann-Kadison-Halmos problem for almost commuting self-adjoint matrices (see [11] and [12]). In that case, the problem has an affirmative solution [20] (see also [8] and [13], and, for a fixed dimension, see [32]). In an infinite dimensional Hilbert space, as observed by Mumford in his new book (p.182), the answer in general is negative (see also Example 4.6 of [8]) in general. The affirmative solution for almost commuting self-adjoint matrices led the further study in the subject of weakly semi-projectivity in -algebras. It also has a deep impact on the Elliott program of classification of amenable -algebras (see [7], [24], [26] and [28]). The Mumford’s problem opens a new door to the future applications of this part of -algebra theory in previous unexpected areas such as DNA and artificial intelligence ([34]).
@@
Let and There exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and are self-adjoint linear operators on with () such that
| (e 1.3) |
@@
Here and are the sets of -synthetic-spectrum, and essential -synthetic-spectrum, respectively (see Definition 2.7). In (e 1.3), is the Hausdorff distance between the two compact subsets of the -dimensional unit cube. In the special case that is full, i.e., it contains the whole then the second part of (e 1.3) is automatically satisfied.
@@
More generally, let be a Hilbert (right) module over a -algebra be the -algebra of all bounded module maps on with adjoints and let be the linear span of rank one module maps of the form (for all ), where Denote by the closure of which is a -algebra. We also have the following:
@@
Let and There exists satisfying the following: Suppose that is a countably generated Hilbert-module over a -unital purely infinite simple -algebra and are self-adjoint bounded module maps with () such that
| (e 1.5) |
@@
(ii) is a countably generated Hilbert module over a -unital purely infinite simple -algebra
@@
-algebra of stable rank one such that has real rank zero and is simple.
@@
The second condition in (e 1.7) means that a Fredholm index vanishes (see Definition 2.4). Note that both case (ii) and case (iii) are generalization of case (i). In case (ii), is stable (whenever is not a finitely generated projective Hilbert module), and is simple (like the Calkin algebra). In case (ii), we keep the feature that is simple and has real rank zero.
Case (i) in Theorem 1.3 is a more definite result than that of Theorem 1.1 in the case We actually prove a slightly more general result (see Theorem 5.3 below). One should also compare (i) in Theorem 1.3 to Theorem 1.1 of [17] (see Remark 5.6). In the case that however, the normal elements may not be approximated by normal elements in with finite spectrum. We also show that condition of the vanishing index associated with the pair of operators is necessary. Since it is well-known that spectrum (and the essential spectrum) of a bounded operator is notoriously unstable regarding perturbation, it is necessary to consider some version of -spectrum (which is usually larger than the spectrum). For we introduce the notion of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebra (see Definition 2.7). We will also show that, unfortunately, the version of Mumford’s problem formulated above has a negative solution in general even the associated Fredholm index vanishes.
In quantum mechanics, one often heard “when the commutators go to zero, we recover the classical system”. If “recover” means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [34] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all show that, mathematically, there are serious (topological) obstacles prevent this type of statements from being true. Nevertheless Theorem 1.1 with some interpretation in Remark 6.6 may provide an alternative statement.
The paper is organized as follows. Section 2 contains some notations used in the paper and gives the definition of -synthetic-spectrum and -near-spectrum for -tuples of self-adjoint operators, and discusses the relationship between them (see also Proposition 2.16). Section 3 is a study of -tuples of almost commuting self-adjoint elements in a unital purely infinite simple -algebra. Section 4 presents the proof of Theorem 1.1 and Theorem 1.2. In Section 5, we will discuss the case of a pair of almost commuting self-adjoint bounded module maps and provide a proof of Theorem 1.3. In section 6, the last section, we show why the Fredholm index does not appear in the statement of Theorem 1.1. We then present an example that an -tuple version of Theorem 1.3 does not hold in general. A final remark was added which contains the justification of the second condition in (e 1.3) and its connection to the Mumford original question.
@@ 2 Preliminaries
Let be a -algebra. Denote by the set of all self-adjoint elements in Denote by the multiplier algebra (the idealizer of in ). If and we write
| (e 2.9) |
For unital -algebra dnote by and the group of invertible elements and the path connected component of containing the identity of respectively.
@@
Let be an infinite dimensional separable Hilbert space. Denote by the -algebra of all bounded linear operators on and by the -algebra of all compact operators on Denote by the quotient map.
@@
Let be a -algebra and be a Hilbert (right) module over (see [15]), or Hilbert -module. Denote by the -algebra of all bounded module maps on with adjoints. A rank one module map is a bounded module map of the form for all (and fixed ), where is the -value inner product on Denote by the linear span of rank one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
@@
Let be a unital -algebra and be an invertible element. Define Note that is a unitary. Define
| (e 2.10) |
Then if and only if when Let be such that is invertible in Then
| (e 2.11) |
@@
Let and In what follows, for each we choose a fixed integer such that Denote We write Then is -dense in Moreover, if
Let be a unital -algebra and be an -tuple of self-adjoint elements in with (). Fix Let be such that if if and is linear in and in Define, for
| (e 2.12) | |||||
| (e 2.13) |
Note that we do not assume that mutually commute and the product in (e 2.13) has a fixed order.
@@
In what follows, we are particularly interested in the case that
@@
Let be a unital -algebra and for some Let us assume that Fix Suppose that is a compact subset and is a unital c.p.c. map such that
| (e 2.19) | |||||
| (e 2.20) | |||||
| (e 2.21) |
@@
We often use the -near spectrum as it associated with a c.p.c. map, whenever the -tuple sufficiently almost commutative as shown in the next proposition (2.11). But it is not unique and depends on the choice of the c.p.c. map. On the other hand, the -synthetic-spectrum is always defined and, in the computation world, it is computable. Their relationship is described in Proposition 2.15 below (see also Proposition 2.16). In practice, it is quite possible that it is more difficult to compute the spectrum of an operator where than to compute the norm of
@@
Fix For any there exists satisfying the following: Suppose that is a unital -algebra and with such that
| (e 2.22) |
@@
Consider and quotient map Let Then Let be the -subalgebra of generated by Then for some non-empty compact subset of Let be the monomorphism given by with
@@
Then, for any and any finite subset there exists satisfying the following: Suppose that is a unital -algebra and are unital c.p.c. maps such that
| (e 2.29) |
@@
| (e 2.30) |
Moreover, if and then, for any and any finite subset we may also require that, for all
| (e 2.31) |
@@
Put Let the quotient map. Then, since is a generating set of by (e 2.33), is a unital homomorphism, and by (e 2.32), Hence
| (e 2.36) |
@@
| (e 2.37) |
This leads a contradiction to (e 2.34). The first part of the lemma then follows.
@@
| (e 2.39) |
Another contradiction. So the lemma follows. ∎
@@
Let be as defined in 2.7, where (). There is satisfying the following: for any unital -algebra and any pair of self-adjoint elements with if then
| (e 2.42) |
@@
Let and Choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebra ), that
| (e 2.43) |
@@
| (e 2.45) |
for any -tuple of self-adjoint elements in with (), by the choice of (and ) and by the second part of (e 2.40) and (e 2.42), we have, for each
| (e 2.46) |
@@
| (e 2.48) |
It follows that (keeping in mind of (e 2.45))
| (e 2.49) |
Combining (e 2.49) with (e 2.47), we obtain
| (e 2.50) |
We conclude that, by (e 2.50) and (e 2.46),
| (e 2.51) | |||
| (e 2.52) | |||
| (e 2.53) |
@@
(1) Suppose that is a unital -algebra and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of and then
| (e 2.55) |
@@
| (e 2.57) | |||
| (e 2.58) | |||
| (e 2.59) |
for any with for all whenever (if ) or (if ). Let We may assume that Then By (e 2.59),
| (e 2.60) |
By (e 2.58) and the choice of applying Lemma 2.12,
| (e 2.61) |
@@
| (e 2.67) |
@@
| (e 2.72) | |||
| (e 2.73) | |||
| (e 2.74) |
for any with if for some Choose We may assume that Hence It follows from (e 2.74) that
| (e 2.75) |
On the other hand, by the choice of (e 2.58), (e 2.73), and applying Lemma 2.12, we have
| (e 2.76) |
@@
Let There exists satisfying the following: Suppose that is a unital -algebra and such that (), and
| (e 2.79) |
@@
Let and be the quotient map. Put Then, by (e 2.80), is normal. Let Then defined by for is a unital monomorphism. By the Choi-Effros Lifting Theorem ([4]), there is a sequence of unital c.p.c. maps such that Let be as in Proposition 2.15. Then, since is a homomorphism, for all large is a -near-spectrum for Applying Proposition 2.15, we may assume that, for all
| (e 2.83) |
Note that, by (e 2.82) and (e 2.83), Let Note that Hence there are such that ()
| (e 2.84) |
@@
| (e 2.85) |
This is a contradiction. So the lemma follows. ∎
@@ 3 Purely infinite simple -algebras
Let be a compact subset of Define to be the union of all even -bricks (in with corners at – see Definition 2.6) which has a non-empty intersection with
@@
Let be a -algebra and a homomorphism. Denote by the restriction.
@@
Let be a compact metric space with property (F). Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism such that
| (e 3.93) |
(see Definition 3.4 for ). Then, for any and any finite subset there exists a set of finitely many mutually orthogonal projections with and such that
| (e 3.94) |
@@
Recall that is finitely generated (). It follows that, for any unital -algebra
| (e 3.95) |
Define by for all Then one computes that, by (e 3.93),
| (e 3.96) |
@@
Recall that, by S. Zhang’s theorem ([39]), has real rank zero. Since is injective, by applying Lemma 4.1 of [7], we obtain a set of mutually orthogonal non-zero projections with such that
| (e 3.99) |
@@
| (e 3.103) |
Define for all Then is a unital homomorphism from to with finite dimensional range and, by (e 3.103), (e 3.100) and (e 3.98), for all
Finally, since in as we arranged earlier, by the definition of and the “Moreover” part of the statement also holds. ∎
@@
Let be a compact subset of Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism. If is any compact subset with property (F) such that where then there is a unital injective homomorphism such that
| (e 3.105) |
@@
| (e 3.106) |
Suppose that is a compact subset of such that Since is compact, there is with such that
| (e 3.107) |
@@
| (e 3.109) |
Applying Theorem 3.5, there is a unital homomorphism with finite dimensional range such that
| (e 3.110) |
@@
| (e 3.115) |
Then, combining with (e 3.110) and (e 3.112), for all
| (e 3.116) | |||||
@@
| (e 3.117) |
Then is a unital injective homomorphism. Moreover, for all by (e 3.109), the definition of above and by (e 3.116),
| (e 3.118) | |||||
To see the second formula in (e 3.105), we may write, from (e 3.117), since factors through
| (e 3.119) |
@@
Suppose that is a compact subset of and suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism. If is any compact subset with property (F) such that then there is a unital injective homomorphism such that
| (e 3.120) |
@@
Let be a compact metric space which is a a compact subset of (for integer ) and be a unital purely infinite simple -algebra. Suppose that is a unital injective homomorphism. Then, for any and any finite subset there exists such that, for any integer with there exists a unital injective homomorphism such that
| (e 3.122) |
@@
Let and There is satisfying the following: If is a unital purely infinite simple -algebra and with () such that
| (e 3.123) |
then there is integer with and and a -brick combination and a unital injective homomorphism such that
| (e 3.124) |
@@
| (e 3.125) |
Let be the commutative -subalgebra generated by Then there exists a unital homomorphism such that Hence for some compact subset of Denote the isomorphism given above. We may assume that
Put Choose any with and By Theorem 3.6, we obtain a -brick combination and a unital injective homomorphism such that
| (e 3.126) |